OpenStudy (lukecrayonz):

Integral? (Antiderivative) of x^(2/5)

OpenStudy (lukecrayonz):

Not sure on the terminology :-)

hartnn (hartnn):

you function to be intrgrated, is of the form x^n , where n = 2/5 so you will use \(\large \int x^ndx = \dfrac{x^{n+1}}{n+1}+c\)

OpenStudy (lukecrayonz):

Sooo x^(2/5)+1=x^7/5/(7/5)+c?

OpenStudy (lukecrayonz):

antiderivative 12x^2-4x+3

hartnn (hartnn):

that 5 x^(7/5)/7 +c was correct :) now we use the fact that integral of sum is sum of integrals. so, \(\int12x^2dx-\int 4xdx +\int3dx \) and constants can be pulled out of the intrgral

hartnn (hartnn):

=\(12\int x^2dx-........\)

hartnn (hartnn):

use same formula of x^n treat 3 as 3x^0 (that is n=0)

OpenStudy (lukecrayonz):

Exactly the same? so just for 12x^2 I'm going to do (12x^3)/3+C and move on?

hartnn (hartnn):

absolutely! :)

OpenStudy (lukecrayonz):

That's pretty easy..

OpenStudy (lukecrayonz):

Okay wait now I have a question. What is the full answer? I just want to see the form you're writing it in

OpenStudy (lukecrayonz):

because 12x^3/3 is 4x^3 and do I use three +C's? so that my final answer would be +3C?

hartnn (hartnn):

yes, most of the integration problems are easy, with few exceptions ofcourse, especially when you have just started it :) with enough practice, you can master the integration thing :) \(\large \dfrac{12x^3}{3}-\dfrac{4x^2}{2}+\dfrac{3x}{1}+c \\ \large = 4x^3-2x^2+3x+c\)

OpenStudy (lukecrayonz):

So only use one +c, so exactly like the other formula just no C until the end

OpenStudy (lukecrayonz):

\[\int\limits_{}^{}\sqrt{x^{11}dx}\]

hartnn (hartnn):

that + c is just there to denote, that there can be a constant after the integration process. you can get the exact value only if you have some initial conditions. so whether you write c or 3c or c1 ....it doesn't matter :)

hartnn (hartnn):

can you write that in the form of x^n ? n=..?

OpenStudy (lukecrayonz):

This one must be different because technically is x^(11)^1/2

hartnn (hartnn):

why would it be different ? isn't it in the same form of x^n ?

OpenStudy (lukecrayonz):

x^11/2?

hartnn (hartnn):

x^11^1/2 can be written as ...? yes, its x^11/2 :) so just put n= 11/2 !

OpenStudy (lukecrayonz):

x^11/2+1=x^13/2/(13/2)+C

hartnn (hartnn):

don't write +1 on left side \(\int x^{11/2}dx = \dfrac{2x^{13/2}}{13}+c\)

OpenStudy (lukecrayonz):

HMM then what about for 7/x-4/x^3?

OpenStudy (lukecrayonz):

If this one isn't different.. lol :(

hartnn (hartnn):

1/x can be written in the form of x^n ?

OpenStudy (lukecrayonz):

x^-1

hartnn (hartnn):

and \(\dfrac{1}{x^3} = x^{???}\) oh and the first part is indeed different

hartnn (hartnn):

\(\huge \int \dfrac{1}{x}dx = \ln |x|+c\) because differentiation of ln x is 1/x

OpenStudy (lukecrayonz):

x^-3

hartnn (hartnn):

correct so for the 2nd part you use x^n formula with n = -3

OpenStudy (lukecrayonz):

7x^-1-4x^-3? Lol :(

hartnn (hartnn):

yes, thats your question you still need to integrate that, right ?

hartnn (hartnn):

for the first part use the new formula i gave for 1/x for the 2nd part , use x^n formula with n=-3

OpenStudy (lukecrayonz):

OOO okay I thought that was the final answer I was like that doesn't seem like much of a change, but okay I get it.

hartnn (hartnn):

yeah so integrate it and tell me what u get ?

OpenStudy (lukecrayonz):

7x^0/0-4x^-2/-2+c

hartnn (hartnn):

why did you use x^n formula for 1/x when i gave you a new formula for it ? the 2nd part "-4x^-2/-2" is correct! just simplify it

hartnn (hartnn):

the reason you cannot use the x^n formula for 1/x is that you get a 0 in the denominator, which is not allowed.

OpenStudy (lukecrayonz):

I don't see why you can use that, you said 1/x so I figured it had to be 1/x to use it?

OpenStudy (lukecrayonz):

\[\frac{ 7 }{ x }-\frac{ 4 }{ x^3 }\]

hartnn (hartnn):

thats the question

OpenStudy (lukecrayonz):

I GOT IT i see

hartnn (hartnn):

so whats your final answer now ?

OpenStudy (lukecrayonz):

Your formula kind of confused me because you put 1/x and I was like, how exactly do you get 1/x? But I see you take the top term and put it on the outside of the antiderivative so it's \[7\int\limits_{}^{}\frac{ 1 }{ x }\]

OpenStudy (lukecrayonz):

7ln(x)+2/x^2+C

hartnn (hartnn):

the top term(7) was a CONSTANT. that is why we can bring it out of the integral if it were variable, then we could have not brought it out of integral

hartnn (hartnn):

and your final answer is correct :)

OpenStudy (lukecrayonz):

Sorry I barely know my terms lol, just know how to do the work :(

OpenStudy (lukecrayonz):

\[4e^{-7x}dx\]

hartnn (hartnn):

ok, since derivative of e^x is just e^x, \(\int e^x dx = e^x+c\)

hartnn (hartnn):

but we have the exponent as -7x here so, we make a substitution

hartnn (hartnn):

let -7x = u we need to find dx now in terms of 'u' and 'du' can you differentiate this ? -7x = u what do u get ?

OpenStudy (lukecrayonz):

Sorry I solved through, -4e^-7x/7+C

hartnn (hartnn):

thats correct, did you understand how to get that ?

OpenStudy (lukecrayonz):

Yes :-) Find f such that f'(x)=5x^2+3x-5 and f(0)=8

hartnn (hartnn):

since we know that integral is anti-derivative the question is just \(\int 5x^2+3x-5 \:\: dx \) right ? and i think you can solve this integral

hartnn (hartnn):

tell me what u get after integrating, then i'll tell you how to get the value of c

OpenStudy (lukecrayonz):

integral: 5x^3/3+3x^2/2-5x+C

hartnn (hartnn):

yes, so your f(x) = 5x^3/3+3x^2/2-5x+c is correct :) now we are given f(0) = 8 so to get f(0), just put x = 0 in f(x) what do u get ?

OpenStudy (lukecrayonz):

8..? lol

hartnn (hartnn):

what u get if you put x= 0 in f(x) = 5x^3/3+3x^2/2-5x+c ?

hartnn (hartnn):

did you mean c= 8 ?

OpenStudy (lukecrayonz):

Yes C=8

hartnn (hartnn):

then you are correct

hartnn (hartnn):

f(x) = 5x^3/3+3x^2/2-5x+8 is the final answer.

OpenStudy (lukecrayonz):

Last question :-)

OpenStudy (lukecrayonz):

\[\int\limits_{}^{}\frac{ u^{5}-2u^{3}+8 }{ 3u }du\]

hartnn (hartnn):

separate out the denominator

hartnn (hartnn):

like (a+b+c)/d = a/d+b/d+c/d

OpenStudy (lukecrayonz):

u^5/3u-2u^3/3u+8/3u

hartnn (hartnn):

yes, so whats u^5/3u = ... ?

OpenStudy (lukecrayonz):

Now it's just applying everything I did before lol

hartnn (hartnn):

yes

OpenStudy (lukecrayonz):

u^3/3-2u-4/u^2+C

hartnn (hartnn):

what have you done :O

OpenStudy (lukecrayonz):

All of it!! :D

hartnn (hartnn):

go term by term :P then you won't do any error u^5/3u = ... ?

OpenStudy (lukecrayonz):

Well it marked right :P

hartnn (hartnn):

LOL! wheen it clearly was incorrect :P

OpenStudy (lukecrayonz):

What?!:( What was wrong? haha

hartnn (hartnn):

i thin i am asking you this for 3rd time u^5/3u =... ?

OpenStudy (lukecrayonz):

u^4/3

OpenStudy (lukecrayonz):

WAIT SO HOW DID I GET IT RIGHT XD

hartnn (hartnn):

thats your 1st term to integrate so, you get (1/3) u^5/5 after integration, right ? i have no idea how you got that right :P

OpenStudy (lukecrayonz):

Let me check my work idk

OpenStudy (lukecrayonz):

Honestly I did an error but MathWay is showing u^3/3 and wolfram is showing something way different.. hrmm lol

OpenStudy (lukecrayonz):

it should be 4.. wth LOL

hartnn (hartnn):

when you can do it on your own, why go to virtual answering sites ? :P

OpenStudy (lukecrayonz):

To check my answer for questions like why this is marked right when its wrong. :O

hartnn (hartnn):

what should be 4 ?

hartnn (hartnn):

= u^5/3u-2u^3/3u+8/3u = u^4/3 -2/3 u^2 +8/3 (1/u) taking the integration = 1/3 (u^5/5) -2/3 (u^3/3) + 8/3 ln |u| +c see if u get this